How do I solve for x with factoring for the following: 25x^2+16=40x.
I'm not sure but i think that I need to move 40x over to the right side so it looks like this:
25x^2+16-40x=0
But I'm unsure.

Answer by John10(297)

Hi,
Good start!
You have 25x^2 - 40 x + 16 = 0
From there, you will factor the left side:
(5x)^2 - 2(5x)(4) + 4^2 = 0
(5x - 4)^2 = 0
You can solve for x from here.
If you have any questions and need some assistance, please email me.
Thank you!